Grade 11Mechanics

a 30g bullets leaves a rifle with a velocity of300 m/s and rifle recoils with a velocity of 0.60 m/s.find mass of rifle

ravandarshan
10 Years agoGrade 11

1 Answer

Amogh Gajare
10 Years ago
This problem can be solved by the principle of conservation of momentum.
By conservation of momentum (P),
          P1 = P2
We know that P = m.v
So, m1v1=m2v2
So substituting values,
0.03 * 300 = 0.6 * m2  as 30 g = 0.03 kg.
Therefore after solving, m2 = 15 kg.